OpenStudy (anonymous):

Help please... Find the intercept: x + y-9= 4

OpenStudy (anonymous):

try getting y by itself and put 0 in for x

OpenStudy (anonymous):

thats your y intercept for x intercept put 0 for y

OpenStudy (anonymous):

(0) + y -9 = 4 +9 +9 y= 13 ?

OpenStudy (anonymous):

x=13-y

OpenStudy (poopsiedoodle):

y=13. Yes. x=13 also. :3

OpenStudy (anonymous):

x+y-9=4 +9 +9 x+y=13 -x -x y=-1x+13

OpenStudy (anonymous):

sorry the intercept is the constant number ( the nmber without a variable)

OpenStudy (anonymous):

Use y=mx+b :)